OpenStudy (nuccioreggie):

Jamie has a deck of 60 sports cards, of which some are baseball cards and some are football cards. Jamie pulls out a card randomly from the deck, records its type, and replaces it in the deck. Jamie has already recorded three baseball cards and nine football cards. Based on these data, what is, most likely, the number of baseball cards in the deck? 12 15 24 30

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

its 15 @Michele_Laino

OpenStudy (michele_laino):

the experimental probability, to get a baseball card, is: favorable outcomes/ possible outcomes= 3/(3+9)=...?

OpenStudy (nuccioreggie):

10

OpenStudy (nuccioreggie):

its 15

OpenStudy (michele_laino):

what is 3/12?

OpenStudy (nuccioreggie):

0.25

OpenStudy (michele_laino):

correct! so the requested number, is given by the subsequent computation: probability* possible outcomes= 0.25*60=...?

OpenStudy (nuccioreggie):

15 i already said that

OpenStudy (michele_laino):

that's right!